OpenStudy (anonymous):

Solve the system using an augmented matrix. Show all work. -x - 3y = -17 2x - 6y = -26 I just can get it ive been researching and watching videos on matrices all day, please help

OpenStudy (abb0t):

\[\left[\begin{matrix}-1 & -3 \\ 2 & -6\end{matrix}\right] = \left(\begin{matrix}-17 \\ -26\end{matrix}\right)\]

OpenStudy (anonymous):

That is how you make it into a matrix right?, but how do you solve the system?

OpenStudy (abb0t):

you want to get it in the form \[\left[\begin{matrix}1 & 0 \\ 0 & 1\end{matrix}\right]\]

OpenStudy (abb0t):

Remember there are three things you can do: 1. Interchange Two Rows 2. Multiply a Row by a Constant 3. Add a Multiple of a Row to Another Row

OpenStudy (abb0t):

You always want a 1 on R1. So next, you can interchange R1 <--> R2

OpenStudy (abb0t):

Do you see the pattern here?

OpenStudy (anonymous):

no not really, im sorry im confused

OpenStudy (abb0t):

Haha, it's ok. These are kind of tough problems.

OpenStudy (abb0t):

You are basically manipulating the matrix to get it in the form as I stated before. That wasy, you only have ONE y = b and ONE x = a.

OpenStudy (anonymous):

okay that makes sense, but how do you get there

OpenStudy (abb0t):

Ok. give me a minute, im doing it right now, and i think I got it right haha. i havent done matrices in a while. but give me a minute to do it. I am only going to do it for the 2x2 tho cuz idk how to do it for a 2x3 on here. but I can draw it 2.

OpenStudy (abb0t):